OpenStudy (anonymous):

A block of mass 4 kilograms, which has an initial speed of 6 meters per second at time t=0, slides on a horizontal surface. b) If a constant friction force of magnitiude 8 newtons is exerted on the block by the surface, determine the following: i. The speed v of the block as a function of time.

OpenStudy (anonymous):

Not 100% sure, but if it helps. When it says, as a function of time. This might mean they don't want you to get an actual numerical answer but produce a "method" for solving it. If I input some time t, I get out the speed (velocity of the block). I'm sure that they have also given you the strength of gravity, maybe 9.8m /s^2. So F = ma, you can get the force acting on the block perpendicular to the horizontal surface. The equations you use here are Newtons equations of motion, this is known as "Kinematics". V = u+ at a = (v-u)/t S = (v+u)/2 * t V^2 = u^2 +2as S = ut + 1/2 at^2 (this is a quadratic) I can't remember about friction forces, but I do remember that every action has an opposite and equal reaction. So the friction must be over combed, and if the force on the block is 8 newtons it will travel at a constant speed, but if you put a little bit more than the 8 newtons you will get it to accelerate. I hope this helps your understanding. Sorry I can't answer it for you.